OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4.

OpenStudy (anonymous):

I don't need the answer..well I do but I would like to be explained how to do it. lol

OpenStudy (asnaseer):

what do you think is the slope of the line y = -3x -4 ?

OpenStudy (anonymous):

umm -3 correct?

OpenStudy (asnaseer):

correct - so if the line you want to find is parallel to y = -3x - 4, then you know it must have the same slope as this. so now you know that the line has slope -3

OpenStudy (asnaseer):

next, do you know the general equation of a straight line?

OpenStudy (anonymous):

not off the top of my head

OpenStudy (asnaseer):

you need to know that in order to solve this question. Look back at your class notes or text book to refresh yourself and then come back to this problem.

OpenStudy (anonymous):

Ok .

OpenStudy (asnaseer):

this web site might also help you: http://www.mathsisfun.com/equation_of_line.html

OpenStudy (anonymous):

y = mx + b?

OpenStudy (anonymous):

Oh! Thank you!

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

But was that right? c;

OpenStudy (asnaseer):

yes - y = mx + b is indeed the general equation of a straight line. here "m" represents the slope of the line and "c" represents where it intersects (or crosses) the y-axis.

OpenStudy (asnaseer):

so in your case, you know that the sope of the line you are looking for is -3, so you know that m=-3, you therefore now know that the equation of the line you want is of the form:\[y=-3x+c\]

OpenStudy (asnaseer):

you are also told that that this line passes through the point (2, 6). which means that you know that when x=2, y=6. You can use this to get:\[y=-3x+c\]therefore:\[6=-3(2)+c=-6+c\]solve this to find "c" and then you can write the final form for the equation of the line being asked for.

OpenStudy (anonymous):

c = 12 ?

OpenStudy (asnaseer):

correct - so you should now be able to write the equation of the line being asked for.

OpenStudy (anonymous):

Oh! I get it now! Thank you!

OpenStudy (asnaseer):

yw - what answer did you get in the end?

OpenStudy (anonymous):

y=-3x + 12

OpenStudy (asnaseer):

perfect! well done! :)

OpenStudy (anonymous):

Thanks! Thank you once again!

OpenStudy (asnaseer):

np - yw :)